# arXiv:2603.01932v1[cs.CV] 02 Mar 2026

### v1

@@ Abstract
Accurate weed mapping in cereal fields requires pixel-level segmentation from UAV imagery that remains reliable across fields, seasons, and illumination. Existing multispectral pipelines often depend on thresholded vegetation indices, which are brittle under radiometric drift and mixed crop–weed pixels, or on single-stream CNN and Transformer backbones that ingest stacked bands and indices, where radiance cues and normalized index cues interfere and reduce sensitivity to small weed clusters embedded in crop canopies. We propose VISA (Vegetation-Index and Spectral Attention), a two-stream segmentation network that decouples these cues and fuses them at native resolution. The radiance stream learns from calibrated five-band reflectance using residual spectral-spatial attention to preserve fine textures and row boundaries that are attenuated by ratio indices. The index stream operates on vegetation-index maps with windowed self-attention to model local structure efficiently, state-space layers to propagate field-scale context without quadratic attention cost, and Slot Attention to form stable region descriptors that improve discrimination of sparse weeds under canopy mixing. To support supervised training and deployment-oriented evaluation, we introduce BAWSeg, a four-year UAV multispectral dataset collected over commercial barley paddocks in Western Australia, providing radiometrically calibrated blue, green, red, red edge, and near-infrared orthomosaics, derived vegetation indices, and dense crop, weed, and other labels with leakage-free block splits. On BAWSeg, VISA achieves 75.6% mIoU and 63.5% weed IoU with 22.8M parameters, outperforming a multispectral SegFormer-B1 baseline by 1.2 mIoU and 1.9 weed IoU. Under cross-plot and cross-year protocols, VISA maintains 71.2% and 69.2% mIoU, respectively. The BAWSeg data, VISA code, and trained models will be released upon publication.
@@ 1 Introduction
To enable supervised learning and deployment-oriented evaluation in cereal systems, we introduce BAWSeg, a four-year UAV multispectral benchmark collected over two commercial barley paddocks near Kondinin, Western Australia (2020–2023) using a DJI Phantom 4 Multispectral platform with RTK positioning and an upward irradiance sensor. Each campaign provides radiometrically calibrated five-band reflectance orthomosaics (blue, green, red, red edge, and near infrared), five derived vegetation indices (NDVI, GNDVI, EVI, SAVI, and MSAVI), and dense pixel annotations for crop, weed, and other classes that explicitly address crop–weed canopy mixing. Unlike existing resources that are either close-range field imagery or aerial datasets with anomaly-style labels, BAWSeg is designed as a reproducible UAV orthomosaic benchmark with consistent radiometry and leakage-free spatial block splits. These splits support three evaluation protocols—within-plot, cross-plot, and cross-year (Fig. 1)—so that generalization across paddocks and seasons can be measured under realistic domain shifts rather than optimistic random sampling.
We evaluate our approach on representative barley paddocks in the Western Australian grainbelt, where herbicide-resistant weeds are prevalent. On within-plot test blocks, VISA achieves a mean Intersection over Union (mIoU) of 0.756 and a weed IoU of 0.635, with 22.8M parameters. Compared with a SegFormer-B1 baseline trained under identical conditions, mIoU increases by about 2% and weed IoU by 3% at similar throughput. Under cross-plot and cross-year evaluations, mean IoUs of 0.71 and 0.69 are maintained, demonstrating robustness under spatial and temporal domain shifts.
@@
  • VISA model: a two-stream network that integrates radiance and vegetation index reasoning for improved segmentation under mixed canopy scenarios.
@@ 2 Related Work
We group prior work into three areas: UAV multispectral weed mapping with index based pipelines, crop–weed semantic segmentation using convolutional encoder–decoder networks on RGB or multispectral inputs, and Transformer or state space models for remote sensing segmentation. These studies provide the baselines and context for our two stream radiance–index model.
@@
UAV multispectral imagery is widely used for site specific weed management and crop–weed discrimination, typically by deriving vegetation indices such as NDVI, GNDVI, SAVI and MSAVI from calibrated red, green, red edge and near infrared bands and then classifying crop, weed and soil with threshold rules or shallow models including support vector machines, random forests and multilayer perceptrons López-Granados (2011); Peña et al. (2013); López-Granados et al. (2016); Wang et al. (2025); De Castro et al. (2018). Object based image analysis combined with random forests has been used to segment homogeneous regions and generate prescription maps in maize and sunflower, while exploiting only local spectral and texture descriptors Castaldi et al. (2017); Pérez-Ortiz et al. (2015); Ciceklidag et al. (2024). Semi supervised systems and feature selection schemes further improve robustness under limited labels by stacking multiple vegetation indices into per pixel or small patch feature vectors for weed mapping in heterogeneous fields Pérez-Ortiz et al. (2016); Tamouridou et al. (2017); Moshou et al. (2017). More recent work integrates multispectral UAV data with ancillary information and ensemble learners to estimate weed density at plot scale in cereal crops Jin et al. (2023); Ibrahim et al. (2025a, b). Across these studies, decisions are usually taken independently for each pixel or patch using hand designed spectral indices and hard labels that assign a single class per pixel, which restricts the use of row level and field level structure and does not explicitly model the uncertainty of mixed pixels where crop plants and weeds occupy the same footprint.
@@
UAV crop–weed mapping has increasingly been formulated as semantic segmentation with convolutional or encoder–decoder backbones such as U-Net, SegNet and DeepLabv3+, adapted from generic vision to field imagery Ronneberger et al. (2015); Badrinarayanan et al. (2017); Chen et al. (2018). Early agricultural deployments trained fully convolutional networks on UAV or ground imagery to delineate crop rows and weed patches but largely followed single-stream designs Lottes et al. (2017); Fawakherji et al. (2023); Shahi et al. (2023). Most studies use RGB or multispectral bands and append a small set of vegetation indices as extra channels, then feed the concatenated tensor into one encoder that must learn spectral contrast and geometric texture without explicit separation of radiance and normalized index cues Lottes et al. (2017); Sa et al. (2018); Fawakherji et al. (2023). Architectures that dedicate an index-aware branch or explicitly decouple spectral ratio information from spatial structure remain uncommon in the UAV literature, as summarized by recent reviews of deep learning for weed detection and drone imaging in agriculture Li et al. (2023); Gupta et al. (2023). Supervision typically relies on hard one-hot pixel labels, so mixed pixels where crop and weed co-occur are forced to a single class and label uncertainty is rarely modeled Li et al. (2023); Shahi et al. (2023). These choices limit robustness under illumination change and can reduce sensitivity to small weed clusters embedded in crop canopies in orthomosaics Sa et al. (2018); Fawakherji et al. (2023); Olsen et al. (2019).
@@
Transformer architectures have been adopted for remote sensing segmentation to capture long range dependencies that are difficult for convolutional encoders. Early work adapts the Vision Transformer and Swin Transformer to aerial and satellite imagery using self attention on patch tokens or shifted windows Dosovitskiy et al. (2021); Liu et al. (2021), followed by hybrids such as SegFormer and related CNN Transformer models with lightweight decoders for high resolution land cover from RGB or multispectral inputs Xie et al. (2021); Zuo et al. (2022). Subsequent methods specialize transformers for remote sensing through tailored tokenization and encoder–decoder designs, including RSTFormer and UNetFormer for urban scenes and river networks and SpectralFormer for hyperspectral land cover classification Li et al. (2022, 2023); Kong et al. (2022). In agriculture, transformer variants are mainly used for crop type mapping and field level weed detection from UAV or satellite imagery with single stream ViT style backbones on reflectance stacks or RGB composites and hard pixel labels Sandoval-Pillajo et al. (2025); Uysal et al. (2024). Recent visual state space models such as VMamba and RS3Mamba extend selective state space layers to semantic segmentation and achieve competitive performance on benchmarks such as LoveDA and ISPRS Vaihingen Guo et al. (2023); Ma et al. (2024), yet most transformer and Mamba based approaches still employ a single feature stream without explicit vegetation index reasoning, structured grouping modules such as Slot Attention or annotation protocols that represent mixed crop–weed canopies and soft label uncertainty.
Figure 2: Planned flight paths for the Kondinin experimental fields. Left, E8 polygon and survey lines planned in GSPro. Right, E2 polygon with the same overlap and altitude parameters. The inset shows a representative interior tile used for quality control of row orientation and texture.
@@ 3 Materials and Methods
This section specifies the complete workflow used to build the BAWSeg benchmark and to develop the VISA segmentation model. It first describes the UAV acquisition setup, radiometric calibration and preprocessing steps that produce georeferenced multispectral orthomosaics, followed by the annotation procedure, patch extraction, and leakage-free block splits that support within-plot, cross-plot, and cross-year evaluation. It then details the two-stream architecture, feature fusion, training objective, and implementation settings used consistently across all experiments.
@@ 3.1 Dataset Construction
We first give details of the BAWSeg dataset since the segmentation model is designed to exploit the information available in this dataset. BAWSeg is a multispectral dataset collected by a UAV from repeated field campaigns. It covers platform and radiometric calibration, mission geometry for uniform coverage, and preprocessing that rectifies, coregisters, orthorectifies, and mosaics all bands. It then outlines patch extraction, leakage-free splits, augmentation, and a dense polygon-based pixel annotation protocol with quality control. The process preserves traceability and consistent radiometry and geometry for reliable training and evaluation.
Figure 3: Example of keypoint inliers for cross band and inter frame alignment. Green circles denote SIFT keypoints and blue lines denote matches retained after RANSAC.
@@ 3.1.1 Acquisition Platform and Configuration
Image capture used a DJI Phantom 4 Multispectral (P4M) RPAS with six synchronized cameras, five narrowband monochrome sensors (blue, green, red, red edge and NIR) and one RGB sensor. An upward irradiance sensor logged downwelling light for every exposure and an RTK module with TimeSync referenced geotags to the camera center. The gimbal was fixed at nadir ($90^{\circ}$). Exposure control operated in manual mode with fixed white balance and ISO for each mission and shutter speed was set to avoid motion blur at the commanded groundspeed. Radiometric reference data were collected with a calibrated Lambertian reflectance panel using nadir frames acquired before and after each mission and whenever the solar zenith changed noticeably during flight, with the irradiance sensor enabled for all captures. Simple capture quality gates monitored shutter speed, saturation, irradiance validity and gimbal status and any pass that violated these checks was repeated at the same altitude and speed. During acquisition, EXIF metadata recorded exposure parameters, RTK position, ellipsoidal height, gimbal yaw, pitch and roll, and irradiance flags. After each mission these fields were parsed into a CSV file with frame identifiers, timestamps, platform pose and quality indicators. All data were stored in a mission specific directory hierarchy indexed by date and site with separate locations for raw multispectral and RGB frames, metadata tables and integrity checksums, which preserves capture conditions for the radiometric normalization and orthomosaic generation described in the next subsection.

### v2

# UAV; multispectral; weed mapping; barley; semantic segmentation; vegetation indices; radiometric calibration; orthomosaic; domain shift; benchmark dataset

@@ Abstract
Accurateweed mapping in cereal fields requires pixel-level segmentation from unmanned aerial vehicle (UAV) imagery that remains reliable across fields, seasons, and illumination. Existing multispectral pipelines often depend on thresholded vegetation indices, which are brittle under radiometric drift and mixed crop–weed pixels, or on single-stream convolutional neural network (CNN) and Transformer backbones that ingest stacked bands and indices, where radiance cues and normalized index cues interfere and reduce sensitivity to small weed clusters embedded in crop canopy. We propose VISA (Vegetation Index and Spectral Attention), a two-stream segmentation network that decouples these cues and fuses them at native resolution. The radiance stream learns from calibrated five-band reflectance using local residual convolutions, channel recalibration, spatial gating, and skip-connected decoding, which preserve fine textures, row boundaries, and small weed structures that are often weakened after ratio-based index compression. The index stream operates on vegetation-index maps with windowed self-attention to model local structure efficiently, state-space layers to propagate field-scale context without quadratic attention cost, and Slot Attention to form stable region descriptors that improve discrimination of sparse weeds under canopy mixing. To support supervised training and deployment-oriented evaluation, we introduce BAWSeg, a four-year UAV multispectral dataset collected over commercial barley paddocks in Western Australia, providing radiometrically calibrated blue, green, red, red edge, and near-infrared orthomosaics, derived vegetation indices, and dense crop, weed, and other labels with leakage-free block splits. On BAWSeg, VISA achieves 75.6% mean Intersection over Union (mIoU) and 63.5% weed Intersection over Union (IoU) with 22.8 M parameters, outperforming a multispectral SegFormer-B1 baseline by 1.2 mIoU and 1.9 weed IoU. Under cross-plot and cross-year protocols, VISA maintains 71.2% and 69.2% mIoU, respectively. The full BAWSeg benchmark dataset, VISA code, trained model weights, and protocol files will be released upon publication.
@@ 1 Introduction
To enable supervised learning and deployment-oriented evaluation in cereal systems, we introduce BAWSeg, a four-year UAV multispectral benchmark collected over two commercial barley paddocks near Kondinin, Western Australia (2020–2023), using a DJI Phantom 4 Multispectral platform with real-time kinematic (RTK) positioning and an upward irradiance sensor. Each campaign provides radiometrically calibrated five-band reflectance orthomosaics (blue, green, red, red edge, and near infrared), five derived vegetation indices, namely normalized difference vegetation index (NDVI), green normalized difference vegetation index (GNDVI), enhanced vegetation index (EVI), soil-adjusted vegetation index (SAVI), and modified soil-adjusted vegetation index (MSAVI) , and dense pixel annotations for crop, weed, and other classes that explicitly address crop–weed canopy mixing. Unlike existing resources that are either close-range field imagery or aerial datasets with anomaly-style labels, BAWSeg is designed as a reproducible UAV orthomosaic benchmark with consistent radiometry and leakage-free spatial block splits. These splits support three evaluation protocols—within-plot, cross-plot, and cross-year (Figure 1)—so that generalization across paddocks and seasons can be measured under realistic domain shifts rather than optimistic random sampling.
We evaluate VISA on BAWSeg under three deployment-oriented protocols, namely within-plot, cross-plot, and cross-year, so that performance can be examined under both in-domain testing and spatiotemporal transfer. The experimental study includes protocol-level evaluation, comparisons with representative baselines, and ablation analysis to assess whether decoupled radiance and vegetation-index modelling improves robustness across paddocks and seasons. Quantitative findings are presented in the Results Section and summarized in the Conclusions Section.
@@
  • VISA model: a two-stream network that integrates radiance and vegetation-index reasoning for improved segmentation under mixed canopy scenarios.
@@ 2 Related Work
We group prior work into three areas: UAV multispectral weed mapping with index-based pipelines, crop–weed semantic segmentation using convolutional encoder–decoder networks on RGB or multispectral inputs, and Transformer or state-space models for remote sensing segmentation. These studies provide the baselines and context for our two-stream radiance–index model.
@@
UAV multispectral imagery is widely used for site-specific weed management and crop–weed discrimination, typically by deriving vegetation indices such as NDVI, GNDVI, SAVI and MSAVI from calibrated red, green, red edge and near-infrared bands and then classifying crop, weed and soil with threshold rules or shallow models including support vector machines, random forests and multilayer perceptrons López-Granados (2011); Peña et al. (2013); López-Granados et al. (2016); Wang et al. (2025); De Castro et al. (2018). Object-based image analysis combined with random forests has been used to segment homogeneous regions and generate prescription maps in maize and sunflower, while exploiting only local spectral and texture descriptors Castaldi et al. (2017); Pérez-Ortiz et al. (2015); Ciceklidag et al. (2024). Semi-supervised systems and feature selection schemes further improve robustness under limited labels by stacking multiple vegetation indices into per-pixel or small-patch feature vectors for weed mapping in heterogeneous fields Pérez-Ortiz et al. (2016); Tamouridou et al. (2017); Moshou et al. (2017). More recent work integrates multispectral UAV data with ancillary information and ensemble learners to estimate weed density at the plot scale in cereal crops Jin et al. (2023); Ibrahim et al. (2025a, b). Across these studies, decisions are usually taken independently for each pixel or patch using hand-designed spectral indices and hard labels that assign a single class per pixel, which restricts the use of row-level and field-level structure and does not explicitly model the uncertainty of mixed pixels where crop plants and weeds occupy the same footprint.
@@
UAV crop–weed mapping has increasingly been formulated as semantic segmentation with convolutional or encoder–decoder backbones such as U-Net, SegNet and DeepLabv3+, adapted from generic vision to field imagery Ronneberger et al. (2015); Badrinarayanan et al. (2017); Chen et al. (2018). Early agricultural deployments trained fully convolutional networks on UAV or ground imagery to delineate crop rows and weed patches but largely followed single-stream designs Lottes et al. (2017); Fawakherji et al. (2023); Shahi et al. (2023). Most studies use RGB or multispectral bands and append a small set of vegetation indices as extra channels, and then feed the concatenated tensor into one encoder that must learn spectral contrast and geometric texture without explicit separation of radiance and normalized index cues Lottes et al. (2017); Sa et al. (2018); Fawakherji et al. (2023). Architectures that dedicate an index-aware branch or explicitly decouple spectral ratio information from spatial structure remain uncommon in the UAV literature, as summarized by recent reviews of deep learning for weed detection and drone imaging in agriculture Li et al. (2023); Gupta et al. (2023). Supervision typically relies on hard one-hot pixel labels, so mixed pixels where crops and weeds co-occur are forced to a single class and label uncertainty is rarely modeled Li et al. (2023); Shahi et al. (2023). These choices limit robustness under illumination change and can reduce sensitivity to small weed clusters embedded in crop canopies in orthomosaics Sa et al. (2018); Fawakherji et al. (2023); Olsen et al. (2019).
@@
Transformer architectures have been adopted for remote sensing segmentation to capture long-range dependencies that are difficult for convolutional encoders. Early work adapts the Vision Transformer and Swin Transformer to aerial and satellite imagery using self-attention on patch tokens or shifted windows Dosovitskiy et al. (2021); Liu et al. (2021), followed by hybrids such as SegFormer and related CNN Transformer models with lightweight decoders for high-resolution land cover from RGB or multispectral inputs Xie et al. (2021); Zuo et al. (2022). Subsequent methods specialize Transformers for remote sensing through tailored tokenization and encoder–decoder designs, including RSTFormer and UNetFormer for urban scenes and river networks and SpectralFormer for hyperspectral land cover classification Liu et al. (2022); Li et al. (2023); Kong et al. (2022). In agriculture, Transformer variants are mainly used for crop type mapping and field-level weed detection from UAV or satellite imagery with single-stream ViT-style backbones on reflectance stacks or RGB composites and hard pixel labels Sandoval-Pillajo et al. (2025); Uysal et al. (2024). Recent Mamba-based segmentation studies have expanded this direction in remote sensing. RS3Mamba introduced a dual-branch visual state-space design for remote sensing image semantic segmentation, while MFMamba extended Mamba to multimodal remote sensing segmentation through image and DSM fusion Ma et al. (2024); Wang et al. (2024). Further 2024–2025 studies explored efficient or multiscale Mamba-based decoders for high-resolution scene parsing, including a semantic-Transformer-assisted real-time model, a Vision Mamba model with multiscale multi-frequency fusion, and LMVMamba with adaptation fine-tuning Ding et al. (2024); Cao et al. (2025); Li et al. (2025). Importantly, for agricultural scenes, EGCM-UNet addressed farmland remote sensing image semantic segmentation, and U-MoEMamba reported Mamba-based UAV segmentation results for cabbage heads under complex field conditions Zheng et al. (2025); Li et al. (2025). Nevertheless, most Transformer- and Mamba-based approaches still employ a single feature stream without explicit vegetation-index reasoning, structured grouping modules such as Slot Attention, or annotation protocols that represent mixed crop–weed canopies and soft label uncertainty.
@@ 3 Materials and Methods
This section specifies the complete workflow used to build the BAWSeg benchmark and develop the VISA segmentation model. It first describes the UAV acquisition setup, radiometric calibration and preprocessing steps that produce georeferenced multispectral orthomosaics, followed by the annotation procedure, patch extraction, and leakage-free block splits that support within-plot, cross-plot, and cross-year evaluation. It then details the two-stream architecture, feature fusion, training objective, and implementation settings used consistently across all experiments.
@@ 3.1 Dataset Construction
We first give details of the BAWSeg dataset, since the segmentation model is designed to exploit the information available in this dataset. BAWSeg is a multispectral dataset collected by a UAV from repeated field campaigns. It covers platform and radiometric calibration, mission geometry for uniform coverage, and preprocessing that rectifies, coregisters, orthorectifies, and mosaics all bands. It then outlines patch extraction, leakage-free splits, augmentation, and a dense polygon-based pixel annotation protocol with quality control. The process preserves traceability and consistent radiometry and geometry for reliable training and evaluation.
@@ 3.1.1 Acquisition Platform and Configuration
Image capture used a DJI Phantom 4 Multispectral (P4M) remotely piloted aircraft system (RPAS) (SZ DJI Technology Co., Ltd., Shenzhen, China) with six synchronized cameras, five narrowband monochrome sensors (blue, green, red, red edge and near-infrared (NIR)) and one RGB sensor. The platform includes an integrated spectral sunlight sensor. In our workflow, the irradiance associated with each single-band image was obtained from the XMP metadata field Irradiance stored in the TIFF output of the aircraft. These per-image irradiance records were retained together with the image metadata, while an RTK module with TimeSync referenced geotags to the camera center. The gimbal was fixed at nadir ($90^{\circ}$). Exposure control operated in manual mode with fixed white balance and ISO for each mission, and shutter speed was set to avoid motion blur at the commanded groundspeed. Radiometric reference data were collected with a calibrated Lambertian reflectance panel using nadir frames acquired before and after each mission and whenever the solar zenith changed noticeably during flight, with the irradiance sensor enabled for all captures. Simple capture quality gates monitored shutter speed, saturation, irradiance validity and gimbal status, and any pass that violated these checks was repeated at the same altitude and speed. During acquisition, EXIF metadata recorded exposure parameters, RTK position, ellipsoidal height, gimbal yaw, pitch and roll, and irradiance flags. After each mission, these fields were parsed into a CSV file with frame identifiers, timestamps, platform pose and quality indicators. All data were stored in a mission-specific directory hierarchy indexed by date and site with separate locations for raw multispectral and RGB frames, metadata tables and integrity checksums, which preserves capture conditions for the radiometric normalization and orthomosaic generation described in the next subsection.
@@ 3.1.2 Flight Mission Geometry and Coverage
Figure 2: Planned Planned flight paths for the Kondinin experimental fields. Left, E8 polygon and survey lines planned in GSPro. Right, E2 polygon with the same overlap and altitude parameters. The parallel survey lines indicate the planned lawnmower flight trajectories within each field polygon. The inset shows a representative interior tile used for quality control of row orientation and texture.
3.1.3 Image Preprocessing
Raw multispectral frames were converted to single-precision arrays, filtered by capture quality flags, and denoised with a $3\times 3$ median kernel applied independently to the R, G, B, NIR and RedEdge bands. We selected this as a mild nonlinear denoising step to suppress isolated impulsive artifacts while preserving stronger local edges before cross-band registration and reflectance processing. However, because the operation is local, very small weed responses near the spatial resolution limit may be weakened when they occupy only one to two pixels. Frames that contained more than 0.5% saturated pixels in any band were discarded, and lens vignetting and distortion were corrected using the vendor calibration so that all subsequent steps operated on rectified images. Cross-band alignment then registered every band to the green reference. SIFT keypoints were extracted per band, descriptor matches were filtered with a Lowe ratio test and RANSAC, and a projective homography was estimated for each band, which is adequate for the near-planar scenes at 120 m altitude. The transform was refined by maximizing the enhanced correlation coefficient on a grid of small windows and registration quality is reported as the median reprojection error, with frames above a 0.5 pixel threshold flagged for reprocessing. Figure 3 illustrates typical inlier correspondences after matching and RANSAC.
Panel-based radiometric calibration converted the registered digital numbers to per-band reflectance that is consistent across sorties. For each frame, the irradiance term was obtained from the per-image XMP metadata field Irradiance recorded by the aircraft. For each band, a scale factor was then derived jointly from the pre-flight and post-flight panel captures and these per-image irradiance values, and the frame reflectance was adjusted for exposure and illumination variation, clipped to a fixed range, and stored as float32. Orthorectification then used RTK geotags and structure from motion in Agisoft Metashape Professional (version 2.0.1, Agisoft LLC, St. Petersburg, Russia) to estimate camera poses and a digital elevation model. All bands of each frame were treated as a rigid group and were projected onto the elevation model into a common ground coordinate system. Seamline optimization and blending within the overlap generated multispectral and RGB orthomosaics without exposure compensation for the analytical bands so that radiometric consistency was preserved. The final per-band mosaics were exported as tiled GeoTIFFs for index computation and patch extraction in the dataset pipeline.
Figure 3: Example of keypoint inliers for cross-band and inter-frame alignment. Green circles denote SIFT keypoints and blue lines denote matches retained after RANSAC.
3.1.4 Annotation and Dataset Finalization
Pixels in the multispectral orthomosaics were labelled at native resolution as barley crop, ryegrass weed, or other, where other covers soil, wheel tracks, and infrastructure. Orthorectified and radiometrically normalized mosaics were tiled, and a vegetation mask derived from NDVI restricted annotation to canopy pixels while nonvegetation and invalid pixels were assigned other or an ignore value. Within this mask, annotators used QGIS (version 3.22 LTR, QGIS Development Team, Open Source Geospatial Foundation (OSGeo), Beaverton, OR, USA) with orthomosaic, NDVI and RGB composites to draw crop and weed polygons, and very small polygons were discarded. The polygons were rasterised by assigning each pixel the class of the polygon that contained its centre, overlaps were resolved in favour of weed polygons so that mixed canopies followed the weed label, and remaining vegetation pixels outside all polygons were kept as other. Each tile was annotated by one operator and checked by a second, and a stratified subset was reannotated to maintain consistent class usage across fields and years. The final labels were stored as single-channel GeoTIFFs with other equal to zero, crop equal to one, weed equal to two and ignore equal to two hundred and fifty five, and during training, this map was converted to one-hot targets $y_{b,c,h,w}\in\{0,1\}$, with the ignore value forming a binary mask that removed those pixels from the supervised loss. Representative examples of the RGB orthomosaic, reflectance map, and final hard label map are shown in Figure 4.